# Thermodynamic coprocessor for linear operations with input-size-independent calculation time based on open quantum system

### v1

@@ Abstract
currents $I_{\kappa,j}(\vec{T})\equiv{J}_{\kappa,j}(\vec{T})/\omega_{\kappa}$. Note, the $\Phi_{j}(T_{j})$ is free from index $\kappa$, and depends only on $T_{j}$.
From Eqs. (22) it is seen that $\Phi_{j}(T_{j})$ depend on $r_{\kappa,j}(\vec{T})$, while $r_{\kappa,j}(\vec{T})$ are free parameters. This set of equations always has a solution. Indeed, let potentials $\Phi_{j}(T_{j})=\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$. Then the resistances required for the decrease in potentials equal $r_{\kappa,j}(\vec{T})=(\Phi_{j}(T_{j})-n_{j}(\omega_{\kappa},T_{j}))/I_{\kappa,j}(\vec{T})$. More over, this is not the unique solution of Eq. (22). For $\Phi_{j}$ such as $\Phi_{j}>\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$, there always exists a set of resistors $r_{\kappa,j}$ that satisfies Eqs. (22). Hence, there are infinitely many sets of $\Phi_{j}$ and $r_{\kappa,j}$ values that satisfy Eqs. (22).
Figure 3: Electrical circuit supporting currents equivalent to the energy flows ${J}_{\kappa,j}(\vec{T})$ in Eq. (1).
Figure 4: Equivalent CS of the OQS. Potentials $\varphi_{\kappa,}$ from Fig. 3 with the same $j$ are connected by common bars through resistors $r_{\kappa,j}$. Potentials $\Phi_{j}$ from Eqs. (22) are applied to these bars, providing the same currents as in Fig. 3.
For the considered computational scheme of vector-matrix multiplication based on OQS in each group of close modes we have $\Phi_{j}(T_{j})=n_{j}(\omega_{\kappa}/T_{j})$ and $r_{\kappa,j}=0$. Also, $\varphi_{0}=0$.
@@
By this analogy, we can expand the class of matrices that can be used in computations similarly to how it is done in MNN. If we need to multiply a matrix $A$ whose rows are normalized vectors with positive and negative elements by a positive vector $\vec{b}$, we can split it into two matrices $A=A_{+}-A_{-}$. The $A_{+}$ cares all positive elements in their places in the matrix $A$ and zeros in other places, while the $A_{-}$ cares all absolute values of negative elements of $A$ in their places in $A$, and zeros in other places. After that, the procedure using Eq. (12) can be applied to both matrices $A_{+}$ and $A_{-}$. The results being subtracted with attenuation deliver $A\vec{b}$.
@@
To summarize, in this paper we show that OQS of bosonic modes, i.e., photons, phonons, magnons, etc., can be used as coprocessors implementing multiple vector-matrix multiplications in a parallel analog way with stochastic matrices. We show that stationary energy flow through one of the OQS’s modes to a reservoir sufficiently colder than others (drain reservoir) is proportional to the scalar product of vectors composed of reservoirs’ occupancies and weighted dissipation rates. We show that this fact and ability to manage dissipation rates of OQS’s modes with close frequencies allow us to implement multiplication of an $n$-dimensional vector by an $m\times n$-dimensional matrix, where $n+1$ is the number of reservoirs and $m$ is the number of the modes having close frequencies. The procedure requires the time needed for the establishment of the stationary energy flows, i.e., the OQS dissipation time, which does not depend on the number of reservoirs (input vector dimension) and the initial state of the OQS. In parallel, vector-matrix multiplications with different matrices are implemented at other close frequency groups of OQS’s modes.
@@
The evolution of OQS under interaction with thermal environments is treated in global approach to dissipation [80, 71, 81] that ensures the fulfillment second law of thermodynamics. Hence, the computations being implemented by energy flows are accompanied by entropy growth due to OQS evolution towards its stationary state. Thus, the presented procedure can be seen as a part of developing fields of thermal linear algebra, thermodynamic computing, and thermal neural networks [77, 31, 78, 79].
@@
Notably, intense interaction with the environment in many applications is considered to be a drawback, i.e., entanglement, coherence, and energy should be preserved as long as possible in the OQS  [124, 95, 125]. In the considered procedure the situation is opposite. The faster OQS dissipates, the faster computations are done.
@@
We develop an electrical analogy for the considered system. We show that the system is equivalent to a CS. The dissipation rates of the OQS multiplied by OQS’s modes’ frequencies can be seen as conductivities, reservoirs’ occupancies as potentials, and stationary energy flows as electric currents.
@@
Also we note, that the similar dependence of energy and particle flows on the weighted occupancy of reservoirs is found in fermionic systems with quadratic Hamiltonians [76], i.e. the same electrical analogy holds true in this case and these systems can also be used for the implementation of vector-matrix multiplication at the nanoscale.

### v2

@@ Abstract
currents $I_{\kappa,j}(\vec{T})\equiv{J}_{\kappa,j}(\vec{T})/\omega_{\kappa}$. Note, the $\Phi_{j}(T_{j})$ is free from index $\kappa$ and depends only on $T_{j}$.
From Eqs. (22) it is seen that $\Phi_{j}(T_{j})$ depend on $r_{\kappa,j}(\vec{T})$, while $r_{\kappa,j}(\vec{T})$ are free parameters. This set of equations always has a solution. Indeed, if potentials $\Phi_{j}(T_{j})=\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$, then the resistances required for the decrease in potentials equal $r_{\kappa,j}(\vec{T})=(\Phi_{j}(T_{j})-n_{j}(\omega_{\kappa},T_{j}))/I_{\kappa,j}(\vec{T})$. More over, this is not the unique solution of Eq. (22). For $\Phi_{j}$ such that $\Phi_{j}>\max_{\kappa}n_{j}(\omega_{\kappa},T_{j})$, there always exists a set of resistors $r_{\kappa,j}$ that satisfies Eqs. (22). Hence, there are infinitely many sets of $\Phi_{j}$ and $r_{\kappa,j}$ values that satisfy Eqs. (22).
Figure 3: Electrical circuit supporting currents equivalent to the energy flows ${J}_{\kappa,j}(\vec{T})$ in Eq. (1). This circuit is the circuit from Fig. 2 repeated for each mode of the OQS.
Figure 4: Equivalent CS of the OQS. Potentials $\varphi_{\kappa,j}$ from Fig. 3 are formed by the connection of the wires with the same $j$ by common bars through resistances $r_{\kappa,j}$. The potentials $\Phi_{j}$ are applied to the bars. The potentials $\Phi_{j}$ and the resistances $r_{\kappa,j}$ are defined by Eqs. (22) and provide the same currents through the wires as in Fig. 3.
For the considered computational scheme of vector-matrix multiplication based on OQS in each group of close modes, we have $\Phi_{j}(T_{j})=n_{j}(\omega_{\kappa}/T_{j})$ and $r_{\kappa,j}=0$.
@@
Using this analogy, we can expand the class of matrices that can be used in computations similarly to how it is done in MNN. If we need to multiply a matrix $A$ whose rows are normalized vectors with positive and negative elements by a positive vector $\vec{b}$, we can split it into two matrices $A=A_{+}-A_{-}$. The matrix $A_{+}$ consists of all positive elements in their places in the matrix $A$ and zeros in other places. The matrix $A_{-}$ consists of all absolute values of negative elements of $A$ in their places in $A$, and zeros in other places. After that, the computations can be done with both matrices $A_{+}$ and $A_{-}$ separately using Eq. (12). The results being subtracted with attenuation deliver $A\vec{b}$.
@@
Conclusion and discussion. In this letter, we show that OQS of bosonic modes, i.e., photons, phonons, magnons, etc., can be used as thermodynamic coprocessors implementing multiple vector-matrix multiplications in a parallel analog way with stochastic matrices. We show that stationary energy flow through one of the OQS’s modes to a reservoir which is sufficiently colder than others (drain reservoir) is proportional to the scalar product of vectors composed of reservoirs’ occupancies and weighted dissipation rates. We show that this fact and ability to manage dissipation rates of OQS’s modes with close frequencies allow us to implement multiplication of an $n$-dimensional vector by an $m\times n$-dimensional matrix, where $n+1$ is the number of reservoirs and $m$ is the number of the modes having close frequencies. The procedure requires the time needed for the establishment of the stationary energy flows, i.e., the OQS dissipation time, which does not depend on the number of the reservoirs (input vector dimension) and the initial state of the OQS. In parallel, vector-matrix multiplications with different matrices are implemented at other close frequency groups of OQS’s modes.
This computational system can be realized by microring resonators in the GHz range (OQS) and heated waveguides (reservoirs). It is expected to reveal a maximal computational rate of $10\div 1000$ TOps/s on the $5\times 5$ cm2 area per a mode of the OQS. However, with nowadays technological opportunities in temperature manipulation, this computational rate reduces to $100$ GOps/s. Nevertheless, even this sufficiently reduced computational rate is comparable with the computational rate of some modern GPU solutions and is of interest in the context of a full-thermodynamic computer development.
@@
The evolution of the OQS under interaction with the thermal environments is treated in the global approach to the dissipation [47, 93, 127] that ensures the fulfillment of the second law of thermodynamics. Hence, the computations being implemented by energy flows are accompanied by entropy growth due to OQS evolution towards its stationary state. Thus, the presented computational scheme can be seen as a part of developing fields of thermal linear algebra, thermodynamic computing, and thermal neural networks [3, 77, 70, 118].
@@
Notably, intense interaction with the environment in many applications is considered to be a drawback, i.e., entanglement, coherence, and energy should be preserved as long as possible in the OQS  [104, 123, 71]. In the considered computational scheme, the situation is opposite. The faster OQS dissipates, the faster computations are done, particularly, the lower the Q-factor of the microring resonator is, the faster computations are done.
@@
We develop an electrical analogy for the discussed computing system. We show that the system is equivalent to a CS. The dissipation rates of the OQS multiplied by the OQS’s modes’ frequencies can be seen as conductivities, reservoirs’ occupancies as potentials, and stationary energy flows as electric currents.
@@
Also, we note that the similar dependence of energy and particle flows on the weighted occupancy of reservoirs is found in fermionic systems with quadratic Hamiltonians [126], i.e., the same electrical analogy holds in this case. Thus, these systems can also be used for the implementation of vector-matrix multiplication at the nanoscale.